OpenStudy (anonymous):

Which of the following tables represents a proportional relationship? A. x 8 9 10 11 y 24 48 72 96 B. x 32 33 34 35 y 96 72 48 24 C. x 32 36 40 44 y 96 112 128 144 D. x 8 16 24 32 y 24 48 72 96

OpenStudy (blank ):

I believe D. This is because: 8/24 = 0.333 16/48 = 0.333 24/72 = 0.333 32/96 = 0.333 They are all the same proportion. :)